OpenStudy (jenniferjuice):

Determine the value of y for the system of equations. x − 3y = 4 2x + y = 8

OpenStudy (***[isuru]***):

x - 3y = 4 ---(1) (1)x 2 , 2x - 6y = 8 ----(2) 2x + y = 8 ----(3) (3) - (2) 2x - 6y - [ 2x + y ] = 8-8 -5y = 0 y = 0